OpenStudy (anonymous):

how many integers between 65/9 and 2024/9 are exactly divisible by 7?

OpenStudy (anonymous):

(65/9) = 7.22 & (2024/9) = 224.8 so the next number be divisible by 7 will be 14 & upper limit will be 224 use A.P for finding the number between nth term (an) = a+(n-1)d i.e 224 = 14 +(n-1)7 i.e n =31

OpenStudy (anonymous):

Here they are {14, 21, 28, 35, 42, 49, 56, 63, 70, 77, 84, 91, 98, 105, 112, 119, 126, 133, 140, 147, 154, 161, 168, 175, 182, 189, 196, 203, 210, 217, 224} There are 31 of them/